gelphielvr:

(algebra 1) how can I determine the growth rate in a y= ab^x

MaxTon:

To find the growth rate, use this formula: Growth rate=b−1 Then convert it to a percentage: Growth rate %=(b−1)×100%

gelphielvr:

@maxton wrote:
To find the growth rate, use this formula: Growth rate=b−1 Then convert it to a percentage: Growth rate %=(b−1)×100%
so is the growth rate b?

MaxTon:

@gelphielvr wrote:
@maxton wrote:
To find the growth rate, use this formula: Growth rate=b−1 Then convert it to a percentage: Growth rate %=(b−1)×100%
so is the growth rate b?
b is not the growth rate, but you use b to calculate it!